OpenStudy (anonymous):

Side length= 9in Perimeter and area of the square

OpenStudy (anonymous):

DONT DELETE UR COMMENTS ON ME XD I WAS LOOKING AND THEN IT WAS GONE!

OpenStudy (anonymous):

Knowing that a square has 4 equal sides you get A=S*S P=S+S+S+S=4s perimeter = 36 area = 81

OpenStudy (anonymous):

The first time i did the area with a side of 4inches which is why i deleted it

OpenStudy (anonymous):

A=BH you can look at it this way